Mr. Barr sold his house for $12,000. How much did the house cost him if his loss was 1/3 of the cost?
Answer by MathGeek87(9)

cost =x
sold by =y
y=x-1/3
y=(2/3)x
x=3*12000/2=3*6000=18000